OpenStudy (anonymous):

can someone explain to me how to find two positive numbers such that the sum of the first and twice the second is 100 and their product is a maximum

OpenStudy (anonymous):

You still there? Do you need help?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

Do you know calculus?

OpenStudy (anonymous):

well haha i'm in the class right now but not really...

OpenStudy (anonymous):

Okay. That's fine.

OpenStudy (anonymous):

well are you going to help?(:

OpenStudy (anonymous):

Do you want just the answer, or how to get the answer?

OpenStudy (anonymous):

how to get it please (:

OpenStudy (anonymous):

how to get it please (:

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

let x= the first number let y= the second number

OpenStudy (anonymous):

so, x+2y=100.

OpenStudy (anonymous):

You are trying to maximize the product yz. so let z= the product = to xy

OpenStudy (anonymous):

Following so far?

OpenStudy (anonymous):

yeah(:

OpenStudy (anonymous):

okay, so solve x+2y=100 for x. x=100-2y z=xy z=(100-2y)y z=y(100-2y) z=100y-2y^2

OpenStudy (anonymous):

Now, just find the vertex of that graph using -b/2a

OpenStudy (anonymous):

Therefore, the first value is 25 and the second value is 50.

OpenStudy (anonymous):

Let me simplify this down a bit, because I was writing out my work which may be a bit confusing.

OpenStudy (anonymous):

So, you have been told that the sum of the first number, twice the second is 100 (I'll use just x and y this time to make things less confusing) so, x+2y=100.

OpenStudy (anonymous):

You are trying to minimize the product, P=xy. So solve the first equation x+2y=100, for x.........x=100-2y. Now substitute this into the product formula ( P=xy)

OpenStudy (anonymous):

ohhh wait i get it! thank you!!!! do you have time to help me with another one?(:

OpenStudy (anonymous):

P = xy = (100 − 2y)y = 100y − 2y^2

OpenStudy (anonymous):

Sure.